### v1

# Proximal Optimal Transport Divergences

Abstract.
We introduce proximal optimal transport divergence, a novel discrepancy measure that interpolates between information divergences and optimal transport distances via an infimal convolution formulation. This divergence provides a principled foundation for optimal transport proximals and proximal optimization methods frequently used in generative modeling. We explore its mathematical properties, including smoothness, boundedness, and computational tractability, and establish connections to primal-dual formulation and adversarial learning. Building on the Benamou-Brenier dynamic formulation of optimal transport cost, we also establish a dynamic formulation for proximal OT divergences. The resulting dynamic formulation is a first order mean-field game whose optimality conditions are governed by a pair of nonlinear partial differential equations, a backward Hamilton-Jacobi and a forward continuity partial differential equations. Our framework generalizes existing approaches while offering new insights and computational tools for generative modeling, distributional optimization, and gradient-based learning in probability spaces.
M.K., L.R-B. and B.Z. are partially funded by AFOSR grant FA9550-21-1-0354. M.K. and L.R-B. are partially funded by NSF DMS-2307115.
1. Introduction
Divergences and metrics are tools for measuring the dissimilarity between probability distributions, enabling the comparison of models with data. They are fundamental in data science and machine learning, where tasks such as sampling from unnormalized distributions or generative modeling depend on selecting a suitable discrepancy measure for learning models. The two primary classes of measures for comparing probability distributions are information divergences and optimal transport distances. Information divergences, such as the Kullback-Leibler (KL) and $f$-divergences, are based on expectations of likelihood-related quantities between distributions. Their ease of computation and optimization from samples makes them widely used in machine learning. However, these divergences are meaningful only when the distributions share the same support. In contrast, optimal transport distances do not require overlapping supports, making them more flexible in certain applications. However, this advantage comes at the cost of significantly higher computational complexity.
In this paper, we introduce the proximal optimal transport (OT) divergence between two probability measures $P$ and $Q$ on some space $X$ and $Y$, denoted by $\mathfrak{D}^{c}_{\varepsilon}$ and defined as an infimal convolution
(1.1) $\mathfrak{D}^{c}_{\varepsilon}(P\|Q)=\inf_{R}\left\{T_{c}(P,R)+\varepsilon\mathfrak{D}(R\|Q)\right\}$
where the infimum is over all probability distributions $R$ on $Y$, $T_{c}$ is the optimal transport cost with some cost function $c:X\times Y\to[0,\infty)$ and $\mathfrak{D}$ is an information divergence. The positive parameter $\varepsilon$ controls the trade-off between the information divergence and the optimal transport cost. Under mild assumptions, this is an information divergence. Infimal convolutions have been previously utilized to define novel forms of divergences based on the dual representation of the $1$-Wasserstein distance. For instance, the recently introduced $(f,\Gamma)$-divergence, [16, 4], combines the dual form of the $1$-Wasserstein distance with $f$-divergences, demonstrating the computational ease and flexibility of $f$-divergences with the advantages of the $1$-Wasserstein distance. This paper extends that concept to encompass general optimal transport distances. Importantly, this work reveals that proximal OT divergence is the foundational mathematical principle behind optimal transport proximals and proximal optimization methods frequently employed in generative modeling. By framing it in this context, we gain deeper insights that allow for the expansion of definitions, properties, and algorithms, encompassing both primal and dual formulations and extending beyond them. Notably, despite its fundamental role, proximal OT divergence has not been explicitly introduced as an independent concept until now.
1.1. Why proximal optimal transport divergences?
Proximal OT divergences possess several appealing mathematical properties that make them particularly useful for comparing data to models or between datasets, even when their supports lie on different manifolds. These properties enable robust and stable learning of target distributions, especially in settings where the data is supported on lower-dimensional structures:
  • Upper bounds imply proximal OT divergences are finite. A key advantage is that they are bounded above by the optimal transport distance and the information divergences, i.e.,
    (1.2) $\mathfrak{D}^{c}_{\varepsilon}(P\|Q)\leq\min\left\{T_{c}(P,Q),\varepsilon\mathfrak{D}(P\|Q)\right\}$
    This ensures finiteness even in the absence of absolute continuity, which enhances their suitability for practical applications. For instance, in generative modeling where $P$ represents the generated distribution and $Q$ is the unknown target distribution accessible only through samples, proximal OT divergence remains finite. In contrast, any information divergence relying on the likelihood ratio $dP/dQ$ may be infinite in such settings as $P$ and $Q$ can be mutually singular.
  • Transport and re-weighting property. The interpolation via the infimal convolution given in (1.1) can be interpreted as a two-step procedure. The first step involves redistributing the probability distribution $P$ into an intermediate probability measure $R$ with an associated cost given by $\mathfrak{D}(R\|Q)$. In the second step, $R$ is mapped to $Q$ via optimal transport, incurring a cost $T_{c}(P,R)$. The optimal intermediate measure denoted by $R^{*}$ that minimizes the total cost and determines the optimal OT divergence (1.1). This property makes them powerful tools for learning manifolds [26], as evidenced by related proximal concepts [4, 5, 37, 54].
  • Smoothness. Another key property is that for uniformly continuous cost functions $c$ and $f$-divergences in (1.1), the variational derivative of $\mathfrak{D}^{c}_{f,\varepsilon}(P\|Q)$ with respect to $P$ remains well-defined without requiring absolutely continuous perturbations relative to $Q$ (e.g. singular ones), which is primarily treated as the target distribution in this paper. This contrasts with classical KL or $f$-divergences, which impose such restrictions. This aspect is critical, as variational derivatives play a central role in machine learning tasks, such as GANs, loss functions, and optimal transport methods, ensuring stability under general conditions.
Transport optimal operator. Under mild assumptions, we establish uniqueness of the optimal intermediate measure $R^{*}$ in (1.1). We formulate the optimal measure as the result of the transport proximal operator on the reference measure, i.e.,
(1.3) $R^{*}=\textbf{prox}^{c}_{\varepsilon\mathfrak{D}(\cdot\|Q)}(P):=\operatorname*{argmin}_{R\in\mathcal{P}(X)}\left\{T_{c}(P,R)+\varepsilon\mathfrak{D}(R\|Q)\right\}.$
which is aligned with the broader concept of proximal operators [35], reinforcing its theoretical significance and practical utility in generative modeling [54]. The uniqueness and the transport proximal operator formulation are particularly relevant for generative modeling, where the objective is to approximate an unknown data distribution $Q$ using only samples. A common approach involves starting with a reference distribution $P$, which is easy to sample from, then optimally transports and reweights it to obtain $R^{*}$ (see second bullet above). When $\varepsilon$ is large, the distribution $R^{*}$ serves as the generative model.
Computational aspects. Our computational tools are built upon primal and dual formulations of proximal OT divergences, with the dual formulation offering a variational representation involving test functions. For example, when the information divergence is the KL-divergence, the dual representation is given by
(1.4) $\mathfrak{D}_{\mathrm{KL},\varepsilon}^{c}(P\|Q)=\sup_{\begin{subarray}{c}(\phi,\psi)\in C_{b}(X\times X),\\ \phi(x)+\psi(y)\leq c(x,y)\end{subarray}}\left\{\mathbb{E}_{P}[\phi]-\varepsilon\ln\mathbb{E}_{Q}[e^{-\psi/\varepsilon}]\right\},$
$C_{b}(X\times Y)$ is the set of continuous and bounded functions on $X\times X$. This representation introduces discriminators, drawing an analogy to adversarial learning methods, [19, 31, 1]. From a computational perspective, the dual formulation enables efficient implementations using convex neural networks [3]. Moreover, the computation of the variational derivative of proximal OT divergences involves the discriminator, defined as the optimal $\phi^{*}$ in (1.4), which allows us to build gradient flows and the associated gradient descent algorithm in probability space [18, 20, 13].
Dynamic formulation of proximal OT divergences. The primal formulation along with the Benamou-Brenier formulation [2], can be interpreted through the lens of Mean-Field Games (MFG) [24], where the MFG framework serves as the Benamou-Brenier analogue for proximal OT divergences. While deeper connections to Hamilton-Jacobi-Bellman (HJB) theory and MFG-based theoretical analysis could provide further insights, we do not explore them here. The MFG formulation enables the use of generative modeling tools [53], to compute the proximal OT divergence and optimal intermediate measure in the primal formulation (1.1), such as generative flows like neural ODEs [11] and OT-flows [33].
1.2. Related work
Special cases of the proximal OT divergences have already been extensively analyzed in both theoretical and computational contexts. In particular, when $T_{c}$ is the $1$-Wasserstein distance (i.e., $c(x,y)=d(x,y)$) and $\mathfrak{D}$ is a $f$ or Rényi divergence, the resulting formulation recovers $(f,\Gamma)$ when $\Gamma$ is the set of all 1-Lipschitz functions, and 1-Wasserstein regularized Rényi divergence introduced in [16, 4] and [7] respectively. In general, these divergences interpolate between integral probability metrics (IPMs) over function spaces $\Gamma$, and $f$ and Rényi divergences. Interesting case is also the maximum mean discrepancy (MMD) regularized $f$ or Rényi divergence where $\Gamma$ is the the unit ball in a reproducing kernel Hilbert space (RKHS), studied in [4] and [7] and early foundational work on the dual representation of MMDs developed in [30]. All these divergences, both in their general form and specific cases, have been extensively used as objective functionals in generative adversarial networks (GANs) and gradient flows, showing their ability for stable training across a broad range of target distributions [6, 20, 7, 18], including learning distributions with heavy tails [13] and a combination of Wasserstein-1 and Wasserstein-2 proximals for manifold learning [21]. Optimal transport proximals have also been instrumental in numerical methods for nonlinear partial differential equations and gradient flows, notably in the celebrated JKO scheme [23], which employs the Wasserstein proximal operator of the KL divergence. For recent advancements in numerical JKO-based schemes, see [10]. More recently, optimal transport proximals have attracted significant interest in generative modeling, particularly in the use of Wasserstein-2 regularized KL divergences for training continuous normalizing flows (CNFs) [17, 34]. Approximations of Wasserstein-2 proximals for both linear and nonlinear energy functionals have also been explored in [25], while score-based generative models have been shown to correspond to the Wasserstein-2 proximal of cross-entropy [54, 53]. As proximal OT divergences can be naturally interpreted as entropic regularization, they share strong connections with Schrödinger bridge problems [32] and various other frameworks based on entropic regularization [12, 36, 14].
Distributionally Robust Optimization (DRO) [39] relies on constructing appropriate ambiguity sets to ensure robustness against distributional uncertainty. Common approaches for modeling misspecification include likelihood-based methods, such as information divergences, and distance-based methods, like the Wasserstein distance. Recent advancements have sought to bridge these two perspectives. A novel method in [8] extends OT-DRO by incorporating conditional moment constraints, while [5] introduces optimal-transport-regularized divergences which have been employed to enhance the adversarial robustness of deep learning models, presenting also applications in privacy and robustness against adversarial attacks. These developments align with special cases of our framework, particularly the finite-dimensional $f$-divergence and KL reformulations. Particular forms of proximal OT divergences have been applied in Bayes-PAC [22], resulting in tractable bounds [50].
2. Proximal optimal transport divergences
We introduce the concept of proximal optimal transport (OT) divergences and establish some key elementary properties that make them particularly useful in computational applications. Building on the Benamou-Brenier dynamic formulation of optimal transport, we also establish a dynamic formulation for proximal OT divergences. The resulting dynamic formulation is a first order mean-field game [24] whose optimality conditions are governed by a pair of nonlinear partial differential equations, a backward Hamilton-Jacobi (HJ) and a forward continuity partial differential equations (PDEs). In this section, we consider general information divergences. In subsequent sections, our results focus specifically on the Kullback-Leibler divergence, though the results extend analogously to any $f$-divergence. All proofs are relegated to the Appendix.
Notation
Let $X$ be a Polish space (i.e., a complete, separable metric space), equipped with Borel $\sigma$-algebra $\mathcal{B}_{X}$ and metric $d:X\times X\to[0,\infty)$. We denote $\mathcal{M}(X)$ to be the set of all finite (signed) Borel measures on $X$ and $\mathcal{P}(X)$ to be the set of all probability measures on $X$. If $X$ and $Y$ are two Polish spaces, and $P\in\mathcal{P}(X)$, $Q\in\mathcal{P}(Y)$, then $\Pi(P,Q)$ is the set of all couplings between $P$ and $Q$, that is, the set of probability measures $\pi\in\mathcal{P}(X\times Y)$ such that $\pi(A\times Y)=P(A)$ and $\pi(X\times B)=Q(B)$ for all Borel sets $A\in\mathcal{B}_{X}$, $B\in\mathcal{B}_{Y}$. We denote $C(X)$ to be the set of continuous functions and $C_{b}(X)$ to be the set of bounded continuous functions on $X$, respectively. For functions $\phi\in C(X)$, $\psi\in C(Y)$, denote $\phi\oplus\psi:=\phi(x)+\psi(y)\in C(X\times Y)$.
2.1. Definition
A proximal OT divergence interpolates between an optimal transport distance and an information divergence via an infimal convolution formulation, preserving the advantageous properties of both. An information divergence $\mathfrak{D}(P\|Q)$ provides a notion of “distance” between probability distribution, satisfying $\mathfrak{D}(P\|Q)\geq 0$ and $\mathfrak{D}(P\|Q)=0$ if and only if $P=Q$. However, they are generally not considered probability metrics. Examples of information divergences include Kullback-Leibler divergence, and more generally, $f$-divergences.
An optimal transport cost, (see e.g. [51, 52, 41]), given a cost function $c:X\times Y\to\mathbb{R}\cup\{+\infty\}$ between $P$ and $Q$ is defined as
(2.1) $T_{c}(P,Q)=\inf_{\pi\in\Pi(P,Q)}E_{\pi}[c],$
where the infimum is taken over all couplings $\pi\in\Pi(P,Q)$, defined as the set of distributions with marginal $P$ and $Q$. Therefore, this formulation seeks the optimal coupling $\pi$ that transports $P$ to $Q$ at minimal cost with respect to $c$. The definition allows for $X$ and $Y$ to be different spaces, though they are often the same. For simplicity, we will assume that $X=Y$ going forward. Let $(X,d)$ be a metric space, then a special class of optimal transport distances is the $p$-Wasserstein distances denoted by $W_{p}(P,Q)$, $p\geq 1$, in which $c(x,y)=d(x,y)^{p}$ (note that when $X=\mathbb{R}^{d}$, $c(x,y)=|x-y|^{p}$), and thus
[…]
Definition 2.1.
Let $c:X\times Y\to[0,\infty)$ be a cost function, and let $\mathfrak{D}(\cdot\|\cdot)$ be an information divergence. The proximal optimal transport divergence of $P\in\mathcal{P}(X)$ with respect to $Q\in\mathcal{P}(Y)$ is defined as the infimum convolution
(2.3) $\mathfrak{D}^{c}_{\varepsilon}(P\|Q)=\inf_{R\in\mathcal{P}(Y)}\left\{T_{c}(P,R)+\varepsilon\mathfrak{D}(R\|Q)\right\}$
For $P,Q\in\mathcal{P}_{p}(X)$ and $T_{c}(P,Q)=W_{p}^{p}(P,Q)$, we use the notation
(2.4) $\mathfrak{D}^{p}_{\varepsilon}(P\|Q)=\inf_{R\in\mathcal{P}(X)}\left\{W_{p}^{p}(P,R)+\varepsilon\mathfrak{D}(R\|Q)\right\}$
Example 2.2.
(Examples of proximal OT divergences)
  1. (1)
    When $T_{c}$ is the Wasserstein-$1$ distance, i.e., $c(x,y)=d(x,y)$ and $\mathfrak{D}$ is a $f$-divergence, the resulting formulation recovers the Lipschitz-regularized $f$-divergences. These divergences have been extensively studied for their significance in the design and theoretical analysis of generative models, particularly in the context of stability and robustness in [4, 16, 20, 13].
  2. (2)
    When $T_{c}$ is the Wasserstein-$p$ distance with $p\geq 2$, and $\mathfrak{D}$ is a KL-divergence, leads to Wasserstein-p regularized KL divergence. In Section 2.4, we derive the dynamic formulation of this specific case of proximal OT divergence through the Benamou-Brenier representation. This dynamic formulation yields optimal trajectories that follow straight-line paths with constant velocity [17]. This property aligns with the geodesic structure of Wasserstein-$p$ space and has been effectively utilized in generative modeling, particularly in OT-flows [34] discussed in Section 5.3.
  3. (3)
    In the recent work [21], a novel approach combines Wasserstein-1 and Wasserstein-2 proximals, leading to a proximal OT divergence where $T_{c}$ corresponds to the Wasserstein-2 distance and $\mathfrak{D}$ is a Lipschitz-regularized $f$-divergence. This formulation ensures the well-posedness of generative flows and enables the robust learning of distributions supported on manifolds.
2.2. Elementary properties of proximal OT divergences
We derive some fundamental properties of proximal OT divergence (2.3), showing its desirable and useful characteristics. To proceed, we impose certain assumptions on the optimal transport cost and the information divergence.
Assumption A-T: The transport cost $T_{c}(P,Q)$ is a divergence. Moreover, the map
(2.5) $(P,Q)\mapsto T_{c}(P,Q)$
is convex and lower semicontinuous in the weak topology (this holds if $c$ is bounded below and lower semi-continuous [41]).
Assumption A-D: The map
(2.6) $(P,Q)\mapsto\mathfrak{D}(P\|Q)$
is jointly convex and lower semicontinuous in the weak topology. Moreover, we assume that the level sets
(2.7) $\{R:\,\mathfrak{D}(R\|Q)\leq\alpha\},$
for any $\alpha\geq 0$, are pre-compact in the weak topology and strictly convex.
These assumptions are satisfied by the KL divergence. Moreover, under these assumptions, the proximal OT divergence is indeed a divergence, inheriting the basic convexity and lower semicontinuity from the original information divergence. In particular, compactness and strict convexity guarantee the existence of a unique minimizer $R^{*}$ in the infimal convolution of the proximal OT divergence. These results are stated in the next two theorems which are proved in Appendix .1.
Theorem 2.3 (Divergence properties).
If the assumptions A-T and A-D are satisfied, then
  1. (1)
    Divergence: $\mathfrak{D}^{c}_{\varepsilon}(P\|Q)$ is a divergence:
    (2.8) $\mathfrak{D}^{c}_{\varepsilon}(P\|Q)\geq 0\quad\text{and}\quad\mathfrak{D}^{c}_{\varepsilon}(P\|Q)=0\iff P=Q$
  2. (2)
    Convexity and lower semicontinuity: The map
    (2.9) $(P,Q)\mapsto\mathfrak{D}^{c}_{\varepsilon}(P\|Q)$
    is convex and lower semi-continuous in the weak topology.
Theorem 2.4 (Proximal properties).
If the assumptions A-T and A-D are satisfied, then
  1. (1)
    Upper bound: The proximal OT divergence is bounded above by the optimal transport distance and the information divergences
    (2.10) $\mathfrak{D}^{c}_{\varepsilon}(P\|Q)\leq\min\left\{T_{c}(P,Q),\varepsilon\mathfrak{D}(P\|Q)\right\}$
  2. (2)
    Proximal minimizer: If $\mathfrak{D}^{c}_{\varepsilon}(P\|Q)<\infty$ then there is a unique minimizer $R^{*}\in\mathcal{P}(Y)$ in (2.3):
    (2.11) $\mathfrak{D}^{c}_{\varepsilon}(P\|Q)=T_{c}(P,R^{*})+\varepsilon\mathfrak{D}(R^{*}\|Q)\,.$
The upper bound (2.10) follows directly by simply setting $R=P$ and $R=Q$ in (2.3). As an important consequence, the proximal divergence is finite whenever either the divergence $\mathfrak{D}(P\|Q)$ or the optimal transport is finite. Notably, this holds without requiring absolute continuity, a crucial distinction from the classical KL divergence, where finiteness typically depends on this condition. For example, when considering the $p$-Wasserstein metrics with the KL-divergence, we observe $\mathfrak{D}^{p}_{\mathrm{KL},\varepsilon}(P\|Q)$ is finite whenever $P$ and $Q$ have $p$-th finite moments without any assumption of absolute continuity of $P$ with respect to $Q$. This property makes proximal OT divergences well-suited for comparing data to models or between datasets, especially when data distributions are concentrated on low-dimensional structures, as suggested by the “manifold hypothesis” in machine learning.
Furthermore, given a target distribution $Q$ and leveraging (2.10), both the ball based on the KL-divergence $B_{\mathrm{KL}}(Q)=\{\mu:\mathfrak{D}_{\mathrm{KL}}(\mu\|Q)\leq 1\}$ and $B_{T_{c}}(Q)=\{\mu:T_{c}(\mu\|Q)\leq 1\}$ based on optimal transport cost, are subsets of the proximal OT divergence ball $B_{\mathrm{ProxOT}}(Q)=\{\mu:\mathfrak{D}^{c}_{\varepsilon}(\mu\|Q)\leq 1\}$. $B_{T_{c}}(Q)$ includes probability distributions that may not be absolutely continuous with respect to $Q$ while $B_{\mathrm{KL}}$ excludes them. However, $B_{T_{c}}(Q)$ may fail to capture heavy-tailed distributions (see Example 1 in [13]), unlike $B_{\mathrm{KL}}$. However, $B_{\mathrm{ProxOT}}(Q)$ includes both such type distributions. This highlights the advantage of proximal OT divergences in enhancing robustness, both in generative modeling tasks and in constructing ambiguity sets for DRO.
The existence of a unique proximal minimizer follows naturally from the compactness of the level sets of the map $R\mapsto\mathfrak{D}(R\|Q)$ in the weak topology. The minimizer $R^{*}$ is crucial for proving many of the results in this paper, and also plays a significant role in applications and algorithms. For example, in generative modeling applications, where $P$ is a reference probability distribution which is easy to sample from and $Q$ is an unknown data distribution accessible only through samples, the probability $R^{*}$ often serves as the generative model produced by the generative algorithm [33, 17, 21]. This insight motivates the definition of a proximal operator that maps the reference measure $P$ to the minimizer $R^{*}$. This transport proximal operator generalizes the proximal operator in $\mathbb{R}^{d}$ [35]. When the transport metric is the Wasserstein distance, it is known as the Wasserstein proximal operator and has been studied in [25].
Definition 2.5 (Transport proximal operator).
Let $Q$ be a fixed probability distribution and assume that the assumptions A-T and A-D are satisfied. We define the transport proximal operator
(2.12) $\textbf{prox}^{c}_{\varepsilon\mathfrak{D}(\cdot\|Q)}:\mathcal{P}(X)\to\mathcal{P}(Y)$
as the mapping that assigns each measure $P\in\mathcal{P}(X)$ to the minimizer
(2.13) $R^{*}=\textbf{prox}^{c}_{\varepsilon\mathfrak{D}(\cdot\|Q)}(P):=\operatorname*{argmin}_{R\in\mathcal{P}(Y)}\left\{T_{c}(P,R)+\varepsilon\mathfrak{D}(R\|Q)\right\}.$
When $T_{c}$ is the Wasserstein-$p$ distance, we denote this operator by $\textbf{prox}^{p}_{\varepsilon\mathfrak{D}(\cdot\|Q)}$ and refer to as the Wasserstein proximal operator. When no confusion arises, we abbreviate the notation to $\textbf{prox}^{c}_{\varepsilon\mathfrak{D}}$ as $Q$ is a fixed probability distribution.
Remark 2.6 (Transport and re-weighting).
The proximal minimizer provides a key insight: the measure $P$ is first (optimally) transported to the intermediate measure $R^{*}$ which is absolutely continuous with respect to $Q$, even if $P$ and $Q$ are not mutually absolutely continuous. Moreover, the support of $R^{*}$ can differ significantly from $P$. The cost paid is the optimal transport cost and essentially depends on the distance over which the mass must be moved. Subsequently, $R^{*}$ is redistributed to $Q$ using the information divergence. In this stage, the support of $R^{*}$ remains contained within the support of $Q$ throughout its construction; any attempt to expand or shift it beyond the support of $Q$ would result in an infinite information divergence. This structural property implies that proximal OT divergences can effectively have the potential to learn manifolds [21]. In particular, some evidence has shown that related proximal divergence frameworks exhibit strong manifold learning capabilities, as demonstrated in [4, 5, 37, 54].
Remark 2.7 (Connection to the JKO scheme).
The Jordan-Kinderlehrer-Otto (JKO) scheme is a time-discretization approach to solving the gradient flow of some energy functional $\mathcal{E}(\rho)$ in the Wasserstein-2 space of probability measures $\mathcal{P}_{2}(X)$. It is closely linked to the transport proximal operator, providing a variational framework that iteratively minimizes a combination of the energy functional and a Wasserstein-2 distance regularization. The continuous-time gradient flow is given by:
(2.14) $\frac{\partial\rho(x,t)}{\partial t}=-\nabla\cdot\left(\rho\nabla\frac{\delta\mathcal{E}(\rho)}{\delta\rho}\right)$
where $\rho(x,t)$ is the probability density evolving over time and $\frac{\delta\mathcal{E}}{\delta\rho}$ is the variational derivative of the energy functional. The right hand side of (2.14) represents the Wasserstein gradient flow. Instead of solving the continuous-time gradient flow equation (2.14) directly, the JKO scheme discretizes time into small intervals and finds an approximate solution at each time-step. Specifically, the scheme with a discrete time step of size $\tau$ is formulated as an iterative minimization
(2.15) $\rho^{*}(\cdot,t_{n+1})=\operatorname*{argmin}_{\rho}\left\{\mathcal{E}(\rho)+\frac{1}{2\tau}W^{2}_{2}(\rho^{*}(\cdot,t_{n}),\rho)\right\}$
The continuous-time limit as $\tau\to 0$ recovers the gradient flow PDE (2.14). Known examples are the heat equation and Fokker-Planck equation which can be seen as the gradient flow of the entropy functional and entropy plus potential energy respectively. In the latter example, by a simple calculation:
[…]
with $f(t)=t\log t$ and $q(x)\propto e^{-V(x)}$, the connections of the JKO with the proximal OT divergences follows by (2.13), i.e.,
$\rho^{*}(\cdot,t_{n+1})=\frac{1}{2\tau}\textbf{prox}^{2}_{2\tau\mathfrak{D_{\mathrm{KL}}}(\cdot\|q)}(\rho^{*}(\cdot,t_{n}))$
starting from some initial probability density $\rho_{0}=\rho$.
2.3. Proximal OT divergences interpolate between optimal transport distances and information divergences
The parameter $\varepsilon$ plays a crucial role in proximal OT divergences. It controls the trade-off between the information divergence term and the optimal transport cost, determining the relative importance of reweighting versus transport. More precisely, $\varepsilon$ scales the information divergence term, affecting how much it contributes to the total cost. For example, in the divergence-dominant regime, i.e., when $\varepsilon$ is large, the information divergence term becomes more significant and even a small value of divergence $\mathfrak{D}(R\|Q)$ contributes significantly to the overall objective as magnified by $\varepsilon$, and in fact, as much as the transport cost. This effect is formalized in the following theorem, demonstrating that the proximal OT divergence is not only a lower bound for both the divergence and the optimal transport (see (2.10)) but also serves as a continuous regularization of both terms.
Theorem 2.8 (Interpolation properties).
Under the assumptions A-T and A-D, the following holds:
  1. (1)
    Divergence regularization by OT. As $\varepsilon\searrow 0$, $\frac{1}{\varepsilon}\mathfrak{D}^{c}_{\varepsilon}$ is increasing and
    (2.16) $\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\mathfrak{D}^{c}_{\varepsilon}(P\|Q)$ $=$ $\mathfrak{D}(P\|Q)$
  2. (2)
    OT regularization by divergence. As $\varepsilon\nearrow\infty$, $\mathfrak{D}^{c}_{\varepsilon}$ is increasing and
    (2.17) $\lim_{\varepsilon\to\infty}\mathfrak{D}^{c}_{\varepsilon}(P\|Q)$ $=$ $T_{c}(P,Q)$
The proof of this theorem is given in Appendix .2 and is essentially the same as the one given in [16, 4] for the family of $(f,\Gamma)$-divergences introduced there.
2.4. Dynamic formulation of proximal OT divergences and mean-field games
Here, we demonstrate that the proximal OT divergences have a continuous-time, flow-based formulation that augments the classical Benamou–Brenier formulation of optimal transport [2] with a mean field game (MFG) structure. Specifically, let $X=Y=\mathbb{R}^{d}$ and consider the transport cost $c(x,y)=\frac{1}{p}\|x-y\|^{p}$ for $p>1$ (the factor $\frac{1}{p}$ is for notational convenience). According to the Benamou-Brenier formula, see e.g [42], optimal transport can be interpreted as the problem of finding an optimal vector field $v_{t}(x)$ that continuously transports the probability measure $P$ into the measure $R$, i.e.,
(2.20) $W_{p}^{p}(P,R)=\min_{\rho,v}\left\{\int_{0}^{1}E_{\rho_{t}}\left[\frac{1}{p}|v_{t}|^{p}\right]dt:\begin{array}[]{c}\partial_{t}\rho_{t}+\nabla\cdot(\rho_{t}v_{t})=0\\ \rho_{0}=P,\,\rho_{1}=R\end{array}\right\}.$
which leads immediately to the following theorem.
Theorem 2.9 (Dynamic formulation of proximal OT divergences).
For $p>1$, the proximal OT divergence $\mathfrak{D}^{p}_{\varepsilon}(P\|Q)$ has the following representation:
$\mathfrak{D}^{p}_{\epsilon}(P\|Q)$ $=\inf_{R\in\mathcal{P}(\mathbb{R}^{d})}\left\{\varepsilon\mathfrak{D}(R\|Q)+W_{p}^{p}(P,R)\right\}$
(2.23) $=\min_{v,\rho}\left\{\varepsilon\mathfrak{D}(\rho_{1}\|Q)+\int_{0}^{1}E_{\rho_{t}}\left[\frac{1}{p}|v_{t}|^{p}\right]dt:\begin{array}[]{c}\partial_{t}\rho_{t}+\nabla\cdot(\rho_{t}v_{t})=0\\ \rho_{0}=P(x)\end{array}\right\}$
The choice of the terminal time chosen to be $1$ in Theorem 2.9 is entirely arbitrary. A more natural choice is $T=\varepsilon$ , and by rescaling time for the flow accordingly, we obtain
(2.26) $\mathfrak{D}^{p}_{T}(P\|Q)$ $=$ $T\inf_{R\in\mathcal{P}(\mathbb{R}^{d})}\left\{\mathfrak{D}(R\|Q)+\frac{1}{T}W_{p}^{p}(P,R)\right\}$
$=$ $T\min_{v,\rho}\left\{\mathfrak{D}(\rho_{T}\|Q)+\frac{1}{T}\int_{0}^{T}E_{\rho_{t}}\left[\frac{1}{p}|v_{t}|^{p}\right]dt:\begin{array}[]{c}\partial_{t}\rho_{t}+\nabla\cdot(\rho_{t}v_{t})=0\\ \rho_{0}=P(x)\end{array}\right\}$
which results in a more convenient form for time-discretization and learning algorithms. We can interpret the parameter $T$ as the time-horizon allotted for the Wasserstein component of the proximal OT divergence to carry out the transport.
Remark 2.10 (Proximal in generative flows).
The dynamic formulation of proximal OT divergence is particularly useful in generative modeling algorithms which learn the optimal vector field $v_{t}^{*}(x)$ and thus the flow driving the source $P$ (often a Gaussian measure) toward the proximal $\rho_{T}^{*}=R^{*}=\textbf{prox}^{c}_{\varepsilon\mathfrak{D}}(P)$. The interpolation properties established in Theorem 2.8 indicate that $\rho_{T}^{*}$ is a good approximation of $Q$, when $T>1$ in (2.26). Specifically, Theorem 2.8 states that $\lim_{T\to\infty}\mathfrak{D}^{c}_{T}(P\|Q)=W_{p}^{p}(P,Q)$. Thus, by the approximation
(2.27) $\mathfrak{D}^{c}_{T}(P\|Q)=T\mathfrak{D}(R^{*}\|Q)+W_{p}^{p}(P,R^{*})\approx W_{p}^{p}(P,Q),\qquad T>1$
$R^{*}$ is a good approximation of $Q$, and hence $\rho^{*}_{T}$ serves as a good surrogate of $Q$.
Remark 2.11 (The mean field-game for proximal OT divergences).
The dynamical formulation of the proximal OT divergence closely resembles the Benamou-Brenier formulation of the Wasserstein transport cost [2]. However, the divergence term and the optimization over the intermediate measure $\rho_{T}=R$ in (2.23) serve as a relaxation of the terminal condition in the Benamou-Brenier formulation. Mathematically, this turns the optimization problem (2.23) into a mean-field game (MFG), [24], with $\mathfrak{D}(\rho_{T}\|Q)$ and $\frac{1}{p}|v|^{p}$ playing the role of the terminal cost and action cost in the MFG formulation respectively. The optimality conditions are a coupled system of non-linear PDEs: a backward Hamilton-Jacobi-Bellman (HJB) equation that governs how an individual agent optimally controls its movement while considering the evolving population distribution $\rho(x,t)$, with a forward continuity equation that describes the dynamics of the density of the agents under a particular velocity field [53]. For example, if the Wasserstein-$2$ distance is chosen in place of the optimal transport cost, the mean-field game yields the following system of coupled PDEs
(2.28) $\begin{dcases}-\frac{\partial U}{\partial t}+\frac{1}{2}|\nabla U|^{2}=0\\ \frac{\partial\rho}{\partial t}-\nabla\cdot(\rho\nabla U)=0\\ U(x,T)=1+\log\frac{\rho(x,T)}{Q(x)},\,\rho(x,0)=P(x).\end{dcases}$
and the optimal velocity field is a time-dependent gradient field $v^{*}(x,t)=-\nabla U(x,t)$.
3. Duality and variational representations for proximal OT divergences
The dual formulation of the proximal OT divergences provides a variational representation of the proximal OT divergences that involves suitable classes of test functions. Variational representations of divergences, integral probability metrics, and optimal transport costs play an important theoretical and practical role. For example these representations are an essential ingredient in generative adversarial networks (GANs) which are constructed as a minmax game between minimization over a paramtetrization of the generator and maximization over the function space (the discriminator step) in the variational representation of the divergence [19]. Here in particular, the dual formulation allows us to compute proximal OT divergences using convex neural networks as discriminators, we refer to Section 5.2 for a complete discussion.
3.1. Duality provides alternative representations of optimal transport distances and information divergences
In this section, we choose the KL-divergence as an information divergence, to present our results. One could also extend it to more general $f$-divergences. We begin by revisiting some well-known facts about optimal transport and KL divergence.
Duality for optimal transport
We briefly summarize some well-known results in optimal transport (see [51, 52, 41] for proofs). Given a cost function $c:X\times Y\to\mathbb{R}\cup\{+\infty\}$ we consider the space
(3.1) $\Phi_{c}=\left\{(\phi,\psi)\in C_{b}(X\times Y),\phi\oplus\psi\leq c\right\}$
(recall $\phi\oplus\psi$ denotes the function $\phi(x)+\psi(y)$). A fundamental result in optimal transport is the following duality result: suppose $c$ is bounded below and lower semicontinuous then we have
(3.2) $T_{c}(P,Q)$ $=\inf_{\pi\in\Pi(P,Q)}E_{\pi}[c]$
$=\sup_{(\phi,\psi)\in\Phi_{c}}\left\{\mathbb{E}_{P}[\phi]+E_{Q}[\psi]\right\}$
Regarding the existence of optimizers, an optimal coupling $\pi^{*}$ always holds given that $X$ and $Y$ are Polish. If $c$ is uniformly continuous and bounded, then there exist optimizers $(\phi^{*},\psi^{*})\in\Phi_{c}$. For general costs functions $c$ we cannot expect the optimizers to be bounded but there exists optimizers $\phi\in L^{1}(P)$, $\psi\in L^{1}(Q)$ satisfying $\phi\oplus\psi\leq c$ provided we assume that
(3.3) $c(x,y)\leq a(x)+b(y)\quad\text{ with }a\in L^{1}(p),b\in L^{1}(Q).$
For example if $c(x,y)=d(x,y)^{p}$, this condition holds for all probability measures in $p$-Wasserstein space defined in (2.2).
KL-divergence
The KL-divergence between two probability measures $R\in\mathcal{P}(Y)$ and $Q\in\mathcal{P}(Y)$ is given by
(3.4) $\mathfrak{D}_{\textrm{KL}}(R\|Q)=\left\{\begin{array}[]{cl}\mathbb{E}_{R}\left[\log\frac{dR}{dQ}\right]&\textrm{ if }R\ll Q\\ +\infty\textrm{ otherwise}\end{array}\right.\,.$
We will use the following variational representation, also known as Donsker-Varadhan formula (see [15] for a proof) which expresses the convex duality between the KL divergence and the cumulant generating function $\Lambda(\psi)=\log\mathbb{E}_{Q}[e^{\psi}]$
[…]
The reader should note that this variational representation implies that we define $\mathfrak{D}_{\textrm{KL}}(R\|Q)$ as +$\infty$+ ∞ if $R$ is not a probability measure. While one could extend $\mathfrak{D}_{\textrm{KL}}(R\|Q)$ to $\mathcal{M}(Y)$ using the definition in the first line (3.4) to obtain a finite value, this leads to another convex duality representation (see e.g. [4]). However, the representation in (3.5) is the one that will be essential for our purposes.
We will also use the (convex) dual formula to (3.5) referred to as the Gibbs variational principle: if $\psi\in C_{b}(Y)$ then
(3.6) $\log\mathbb{E}_{Q}[e^{\psi}]=\sup_{R\in\mathcal{P}(Y)}\left\{\mathbb{E}_{R}[\psi]-\mathfrak{D}_{\textrm{KL}}(R\|Q)\right\},$
and the supremum is attained when $R$ is such that
(3.7) $\frac{dR}{dQ}=\frac{e^{\psi}}{\mathbb{E}_{Q}[e^{\psi}]}\,.$
3.2. Duality for proximal OT divergences via variational representations
In this section, we establish a dual formulation for the proximal OT divergence when KL divergence is combined with an optimal transport cost. This is achieved by leveraging the variational representations given in (3.6) and (3.2) respectively. In particular, the infimum convolution formula yields the following duality result.
Theorem 3.1.
(Duality for proximal OT-divergence) Suppose $X$ and $Y$ are Polish spaces and that $c$ is lower-semicontinuous and bounded below. Given $P\in\mathcal{P}(X)$ and $Q\in\mathcal{P}(Y)$ suppose that $c(x,y)\leq a(x)+b(y)$ with $\mathbb{E}_{P}[a]<\infty$ and $\mathbb{E}_{Q}[b]<\infty$. Then we have the following variational representations
(3.8) $\mathfrak{D}_{\mathrm{KL},\varepsilon}^{c}(P\|Q)$ $=\inf_{R\in\mathcal{P}(Y)}T_{c}(P,R)+\varepsilon\mathfrak{D}_{\mathrm{KL}}(R\|Q)$
$=\sup_{(\phi,\psi)\in\Phi_{c}}\left\{\mathbb{E}_{P}[\phi]-\varepsilon\log\mathbb{E}_{Q}[e^{-\psi/\varepsilon}]\right\}{}$
The proof of Theorem 3.1, at this level of generality, is rather long and is given in Section .3. The general idea is to use a convex analysis argument in the case where $X$ and $Y$ are compact and $c$ is continuous and then prove the general case using several approximation arguments and lower semicontinuity. In this respect, the strategy of the proof follows quite closely the proof of duality for optimal transport given in [51] and use many of the technical estimates and ideas in there.
For now, we are content with a formal proof based on the dual formulation of the transport cost $T_{c}(P,R)$ and the Gibbs variational principle for the KL-divergence. Shamelessly by exchanging infimum and supremum–the key technical step, which could likely be justified for compact spaces using an appropriate minimax theorem– we obtain
$\mathfrak{D}_{\mathrm{KL},\varepsilon}^{c}(P\|Q)$ $=\inf_{R\in\mathcal{P}(Y)}\left\{T_{c}(P,R)+\varepsilon\mathfrak{D}_{\textrm{KL}}(R\|Q)\right\}$
[…]
$=\sup_{(\phi,\psi)\in\Phi_{c}}\inf_{R\in\mathcal{P}(Y)}\left\{\mathbb{E}_{P}[\phi]+E_{R}[\psi]+\varepsilon\mathfrak{D}_{\textrm{KL}}(R\|Q)\right\}\quad(\inf\leftrightarrow\sup)$
$=\sup_{(\phi,\psi)\in\Phi_{c}}\left\{\mathbb{E}_{P}[\phi]-\varepsilon\sup_{R\in\mathcal{P}(Y)}\left\{E_{R}[-\psi/\varepsilon]-\mathfrak{D}_{\textrm{KL}}(R\|Q)\right\}\right\}$
[…]
Example 3.2.
(Lipschitz-regularized $f$-divergences) In Example 2.2, the Lipschitz-regularized $f$-divergences [20] arise as a special case of proximal OT divergences, where the information divergence is given by an $f$-divergence and Wasserstein-1 is the optimal transport cost part. It can be readily shown that a $c$-concave function, defined in Appendix .3, corresponds exactly to a 1-Lipschitz function, and vice versa. Moreover, in this case, we have $\psi=\phi^{c}=-\phi$ (see Lemma .1) and thus
$D_{\mathrm{KL}}^{{\mathrm{Lip}}_{\frac{1}{\varepsilon}}(X)}(P\|Q)=\sup_{\phi\in{\mathrm{Lip}}_{\frac{1}{\varepsilon}}(X)}\left\{\mathbb{E}_{P}[\phi]-\ln\mathbb{E}_{Q}[e^{\phi}]\right\}=\frac{1}{\varepsilon}\mathfrak{D}_{\mathrm{KL},\varepsilon}^{1}(P\|Q)$
where ${\mathrm{Lip}}_{\frac{1}{\varepsilon}}(X)$ is the set of Lipschitz functions on $X$ with Lipschtz constant $1/\varepsilon$. This can be extended to the dual form of the Lipschitz-regularized $f$-divergences as $D_{f}^{{\mathrm{Lip}}_{\frac{1}{\varepsilon}}(X)}(P\|Q)=\frac{1}{\varepsilon}\mathfrak{D}_{f,\varepsilon}^{1}(P\|Q)$.
Next, we derive the existence and characterization of the solution $R^{*}$ to the infimal convolution problem (2.3) as well as the optimizers $(\phi^{*},\psi^{*})$ in the dual representation (3.8). Furthermore, the three optimizers $R^{*},\phi^{*},\psi^{*}$ are related in such a way that they reinforce the transport and re-weighting interpretation of proximal OT divergences as discussed in Remark 2.6 and stated in Theorem 3.3 (d) and further explained thereafter.
Theorem 3.3.
(Optimizers for the proximal OT divergence) Assume that $X$ and $Y$ are Polish spaces and that $c$ is lower-semicontinuous and bounded below. Let $P\in\mathcal{P}(X)$ and $Q\in\mathcal{P}(Y)$ and assume that $c(x,y)\leq a(x)+b(y)$ with $\mathbb{E}_{P}[a]<\infty$ and $\mathbb{E}_{Q}[b]<\infty$.
  • $(a)$
    There exists a unique minimizer $R^{*}$ in (2.3).
  • $(b)$
    If $c$ is bounded and uniformly continuous then there exists a maximizer pair $(\phi^{*},\psi^{*})\in\Phi_{c}$ given in (3.8). Up to a constant shift $(\phi^{*},\psi^{*})\to(\phi^{*}-a,\psi^{*}+a)$, $\phi^{*}$ is unique $P$ almost surely and $\psi$ is unique $Q$ almost surely.
  • $(c)$
    In general, the supremum is attained for a pair $(\phi^{*},\psi^{*})\in L^{1}(P)\times L^{1}(Q)$ with the same uniqueness property as for $c$ continuous.
  • $(d)$
    The optimizer $R^{*}$, $\phi^{*},\psi^{*}$ are related in the following manner:
    1. (1)
      The pair $(\phi^{*},\psi^{*})$ is the optimizer in the dual representation of the $T_{c}(P,R^{*})$, i.e.
      (3.9) $T_{c}(P,R^{*})=\mathbb{E}_{P}[\phi^{*}]+E_{R^{*}}[\psi^{*}]\,.$
    2. (2)
      The Radon-Nikodym derivative of $R^{*}$ with respect to $Q$ is given by
      (3.10) $\frac{dR^{*}}{dQ}=\frac{e^{-\psi^{*}/\varepsilon}}{\mathbb{E}_{Q}[e^{-\psi^{*}/\varepsilon}]}$
Note that all optimizers $R^{*}$, $\phi^{*},\psi^{*}$ depend on $\varepsilon$; however, for notational simplicity, we omit this dependence throughout. The proof of this theorem is given in Section .3. The technical part consists in proving the existence of optimizers. Here, we focus on explaining the relationship between the optimizers. Assuming we have established both the duality and the existence of maximizers $(\phi^{*},\psi^{*})$ (.24), then we have
(3.11) $E_{\pi^{*}}[c]+\varepsilon\mathfrak{}R^{*}\|Q)=\mathbb{E}_{P}[\phi^{*}]-\varepsilon\ln\mathbb{E}_{Q}[e^{-\psi^{*}/\varepsilon}]$
where $\pi^{*}$ is the optimal transport plan between $P$ and $R^{*}$. Equivalently
(3.12) $\underbrace{E_{\pi^{*}}[c-\phi^{*}-\psi^{*}]}_{\geq 0}+\varepsilon\underbrace{\left\{\mathfrak{}R^{*}\|Q)-E_{R}[-\psi^{*}/\varepsilon]+\log\mathbb{E}_{Q}[e^{-\psi^{*}/\varepsilon}]\right\}}_{\geq 0}=0$
and, since both terms must vanish, we get exactly the characterizations of the optimizers in (3.9) and (3.10) in Theorem 3.3.
Part $(d)$ in Theorem 3.3 can be interpreted as a transport/re-weighting process (see also Remark 2.6) where (3.9) is the transport of $P$ to $R^{*}$ and the emerging Gibbs distribution in (3.10) describes the re-weighting of $R^{*}$ to $Q$ with the weighting determined by $\psi^{*}/\varepsilon$. In generative modeling, this process enables the generation of new data by transporting and re-weighting samples. At this stage, we recall the Monge formulation and establish a connection between the optimal transport map and the gradient of $\phi^{*}$ given in Part $(d)$ of Theorem 3.3. Let $P,Q\in\mathcal{P}(\mathbb{R}^{d})$. According to the Monge formulation, we have that the optimal transport problem is to find a measurable map $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ that pushes $P$ onto $Q$, minimizing the optimal transport cost:
[…]
subject to $T_{\#}P=Q$. The optimal coupling $\pi^{*}\in\Pi(P,Q)$, see (2.1), is then written as $\pi^{*}=(\mathrm{Id},T^{*})_{\#}P$. For the special case $c(x,y)=\frac{1}{2}\|x-y\|^{2}$, the corresponding optimizer $\phi^{*}$ given in Part $(d)$ of Theorem 3.3 is linked to the optimal transport map as follows: $T^{*}(x)=x-\nabla\phi^{*}(x)$ and $T^{*}(x)=\nabla u$ a.e., where $u(x)=\frac{\|x\|^{2}}{2}-\phi^{*}(x)$, see Theorem 4.2 in [44].
Remark 3.4.
(Alternative expression for the solution of JKO scheme at each step) The iterative minimization in the JKO scheme can be expressed using the transport proximal operator:
$\rho^{*}(\cdot,t_{n+1})=\frac{1}{2\tau}\textbf{prox}^{2}_{2\tau\mathfrak{D_{\mathrm{KL}}}(\cdot\|q)}(\rho^{*}(\cdot,t_{n})),$
see Remark 2.7. From Theorem 3.3(d)2), the corresponding proximal minimizer satisfies
(3.14) $\frac{\rho^{*}(\cdot,t_{n})}{q}=\frac{e^{-\psi_{n}^{*}/2\tau}}{2\tau\mathbb{E}_{Q}[e^{-\psi_{n}^{*}/2\tau}]}\textrm{\;\;which implies \;}\rho^{*}(\cdot,t_{n})\propto e^{-\psi_{n}^{*}/2\tau-V}\,.$
This expression aligns with Proposition 8.7 in [43].
3.3. Applications of duality: data processing inequalities for proximal OT divergences
In this section, we first show that the proximal $p$-Wasserstein divergence is invariant under metric isomorphism. Next, we establish a version of the data-processing inequality,[49]. Finally, leveraging the duality formula, we prove that the proximal OT divergence with $p$-Wasserstein $\mathfrak{D}_{p}^{\varepsilon}$ is additive for product measures.
Theorem 3.5.
(Invariance under isomorphisms) Suppose $T$ is a metric isomorphism of $X$ (i.e. $T:X\to X$ is invertible and preserve distance, $d(Tx,Ty)=d(x,y)$ for all $x,y$). Then
(3.15) $\mathfrak{D}^{p}_{\varepsilon}(T_{\#}P\|T_{\#}Q)=\mathfrak{D}^{p}_{\varepsilon}(P\|Q)$
[…]
Theorem 3.6.
(Data processing inequality) Suppose $X=Y$ and $K$ is a kernel from $X$ to $X^{\prime}$ which maps $C_{b}(X^{\prime})$ into $C_{b}(X)$. Then we have
(3.16) $\mathfrak{D}^{c}_{\varepsilon}(K_{\#}P\|K_{\#}Q)\leq\mathfrak{D}^{Kc}_{\varepsilon}(P\|Q)$
where $Kc$ is the cost function on $C$ defined by
(3.17) $Kc(x,y)=\int_{X}\int_{Y}c(x,y)K(x,dx)K(y,dy)$
Theorem 3.7.
(Additivity for product measures) Assume $X=Y=\mathbb{R}^{n}$ and $c(x-y)=\|x-y\|^{p}$. If $P=P_{1}\times P_{2}$ and $Q=Q_{1}\times Q_{2}$ are product measures with $P_{1},Q_{1}\in\mathcal{P}(\mathbb{R}^{n_{1}})$ and $P_{2},Q_{2}\in\mathcal{P}(\mathbb{R}^{n_{2}})$ then the proximal OT is additive
(3.18) $\mathfrak{D}_{\mathrm{KL},\varepsilon}^{p}(P_{1}\times P_{2}\|Q_{1}\times Q_{2})=\mathfrak{D}_{\mathrm{KL},\varepsilon}^{p}(P_{1}\|Q_{1})+\mathfrak{D}_{\mathrm{KL},\varepsilon}^{p}(P_{2}\|Q_{2})$
The proof relies on both the primal and dual representations of the divergence. By restricting the infimum to product intermediate measures, we obtain one side of the inequality in (3.18), while restricting the supremum to functions of the form $\phi(x)=\phi_{1}(x_{1})+\phi_{2}(x_{2})$ (and similarly for $\psi$) yields the reverse inequality. The full details are provided in Section .4.
4. The first variation of the proximal OT divergence
In this section we show that the variational derivative of the proximal OT divergence around a fixed measure $Q$ exists under very broad conditions and is practically computable. The existence and the evaluation of variational derivatives of proximal OT divergences involves the discriminator (i.e., the optimal $\phi^{*}$ in (3.8)) in the dual variational representation of proximal OT divergences. Here we show that the variational derivative is always well-defined for any probability measure perturbations for which the proximal OT divergence remains finite. We note that such perturbations can be singular and do not necessarily need to be absolutely continuous with respect to $Q$ where the variational derivative is calculated. Therefore, variational derivatives of proximal OT divergences exist in “all directions”, not just along absolutely continuous perturbations of the baseline $Q$ directions, as is for instance in KL and other $f$-divergences. A key practical outcome of this observation is that machine learning algorithms that involve the optimization of $f$-divergences of any type, will be provably stabilized when a proximal OT divergence is deployed instead of an $f$-divergence. For instance, in gradient-based algorithms used to solve the optimization problem $\min_{\theta}\mathfrak{D}_{\mathrm{KL},\varepsilon}^{c}(P^{\theta}\|Q)$, it is necessary to compute the gradient
[…]
involves the variation derivative of $\frac{\delta\mathfrak{D}_{\mathrm{KL},\varepsilon}^{c}(P\|Q)}{\delta P}$ which must be well-defined for the gradient to be properly evaluated. Moreover, the construction of gradient flows in the space of probability measures also relies critically on this variational derivative as discussed in detail in Section 6.
4.1. Proximal OT divergences have well-defined variational derivative
From Theorem 3.3, the optimizers $(\phi^{*},\psi^{*})$ are uniquely determined in $\textrm{supp}(P)$ and $\textrm{supp}(Q)$ respectively, up to additive constant, i.e., that is we can replace $(\phi^{*},\psi^{*})$ by $(\phi^{*}+a,\psi^{*}-a)$ for any $a\in\mathbb{R}$. To compute the directional derivative of the optimal transport divergence, we extend $\phi^{*}$ to all of $X$. This extension follows a standard approach repeatedly used in optimal transport theory, i.e., we define
(4.1) $\widehat{\phi}^{*}(x)=\inf_{y\in\textrm{supp}(Q)}\left\{c(x,y)-\psi^{*}(y)\right\}\,.$
Clearly $\widehat{\phi}^{*}(x)=\phi^{*}(x)$ for all $x\in\textrm{supp}(P)$ and by definition $\widehat{\phi}^{*}(x)$ is the largest function $h(x)$ such that $h(x)=\phi^{*}(x)$ on $\textrm{supp}(P)$ and $h(x)+\psi^{*}(y)\leq c(x,y)$ for $x\in X$ and $y\in\textrm{supp}(Q)$. Our next theorem states that the variational derivative is equal to $\widehat{\phi}^{*}(x)$ and extend the results proved in [16, 20] for the proximal 1-Wasserstein divergence. For simplicity, we focus on the case where $c$ is bounded and uniformly continuous, though the result likely extends to more general cost functions as well.
Theorem 4.1.
(Variational derivative of the proximal OT-divergence) Suppose $c$ is uniformly continuous. Let $\rho\in\mathcal{M}(X)$ be a finite signed measure with total mass 00, $\rho(X)=0$, such that $\rho=\rho_{+}-\rho_{-}$ and $\rho_{\pm}\in\mathcal{P}(X)$ are mutually singular. Assume that for some $\alpha_{0}>0$, $P+\alpha\rho\in\mathcal{P}(X)$ if $\alpha\leq\alpha_{0}$. Then we have
[…]
where $\widehat{\phi}^{*}$ is given by equation 4.1 and $(\phi^{*},\psi^{*})$ are the optimizer in 3.8.
The proof of the theorem is given in Appendix .5. In the sequel, we will use the following symbolic notation involving the variational derivative,
(4.3) $\frac{\delta\mathfrak{D}_{\mathrm{KL},\varepsilon}^{c}(P\|Q)}{\delta P}=\widehat{\phi}^{*}.$
This allows to express Theorem 4.1 in a concise form.
Remark 4.2.
From this theorem (applied to $\rho$ and $-\rho$), we conclude that if $P+\alpha\rho\in\mathcal{P}(X)$ for $\alpha$ in a neighborhood of 00 then $\mathfrak{D}_{\mathrm{KL},\varepsilon}^{c}(P+\alpha\rho\|Q)$ is actually differentiable at 00. However, for differentiability to hold, $\rho$ must be absolutely continuous with respect to $P$. In the general case, we only obtain a one-sided derivative, but this is precisely what is required for constructing a gradient descent algorithm in probability space.
5. Algorithmic aspects of proximal OT divergences
The computational aspects of proximal divergences can be broadly categorized into two interrelated challenges: (1) computing or estimating the proximal OT divergence itself using its primal and dual formulation, and (2) computing the proximal operator $\textbf{prox}^{c}_{\varepsilon\mathfrak{D}}$ given in Definition 2.5. Notably, these are computable formulations of the proximal OT divergence and its associated operators, for which various numerical methods have been proposed in the literature (see, e.g., ).
5.1. Proximal OT divergences for discrete distributions
Proximal OT divergences are computationally tractable due to their dual and dynamic formulations, which offer two alternative algorithmic, as we discuss in the following sections. We first focus on computing the proximal OT divergence in the discrete setting, particularly when $P$ and $Q$ are empirical measures. In this case, the dual variational formulation of the proximal OT divergence directly translates into a convex optimization problem, allowing for efficient numerical computation. Specifically, let $P=\sum_{i=1}^{N}a_{i}\delta_{x^{i}},Q=\sum_{j=1}^{M}b_{j}\delta_{y^{i}}$ be two empirical measures where $(a_{i})_{i=1}^{N},(b_{j})_{j=1}^{M}$ are weights in the probability simplex. We write the intermediate measure as $R=\sum_{i=1}^{N}w_{i}\delta_{y_{j}}$ ensuring that $R$ remains absolutely continuous with respect to $Q$. The primal formulation of the proximal OT divergence with the KL divergence is given by
(5.1) $\mathfrak{D}^{c}_{\varepsilon}(P\|Q)=\min_{(w_{i})}\left\{\min_{\pi}\langle C,\pi\rangle+\epsilon\sum_{i=1}^{M}w_{i}\log\frac{w_{i}}{b_{i}}\right\},$
where $C\in\mathbb{R}^{N\times M}$ is the matrix with entries $C_{i,j}=c(x^{i},y^{j})$ and $\pi\in\mathbb{R}^{N\times M}$ is a (discrete) transport plan from $P$ to $Q$, i.e., $\pi\in\Pi(a,c)$ where
(5.2) $\Pi(a,c)=\left\{\pi\in\mathbb{R}^{N\times M},\;\pi\geq 0,\;\pi 1_{N}=a,\;\pi^{T}1_{M}=c\right\}.$
Alternatively, the proximal OT-divergence can be written using the dual formulation in terms of potential $\phi\in\mathbb{R}^{N},\psi\in\mathbb{R}^{M}$ as
(5.3) $\mathfrak{D}^{c}_{\varepsilon}(P\|Q)=\max_{\phi\oplus\psi\leq C}\left\{\sum_{i=1}^{N}\phi_{i}a_{i}-\epsilon\log\sum_{j=1}^{M}b_{j}e^{-\psi_{j}/\epsilon}\right\},$
where the constraint means that $\phi_{i}+\psi_{j}\leq C_{ij}$ for $i=1,\dots,N$ and $j=1,\dots,M$. Moreover, the weights $w_{i}$ of $R$ are given by
(5.4) $w_{j}=\frac{e^{-\psi_{j}/\varepsilon}}{\sum_{l}e^{-\psi_{l}/\varepsilon}}$
which comes from (3.10). The problem in (5.3) has a concave objective (by noting that the log-sum-exp function is convex) with linear inequality constraints. The problem can be solved using convex programming tools.
5.2. Neural network approximations to proximal OT divergences
Let us consider the proximal OT divergence with the KL divergence and 2-Wasserstein distance, i.e., the objective for the quadratic cost $c(x,y)=\frac{1}{2}\|x-y\|^{2}$. The dual representation (3.8) for proximal OT divergences requires optimizing over the set $\Phi_{c}$ which poses significant computational challenges. To address this, we use the convexification trick introduced in [51]. Building on the approach in [27], we reformulate the problem by enforcing convexity constraints on the function space. This reformulation allows us to express proximal OT divergences as a minimax optimization over convex functions, enabling the use of input convex neural networks (ICNNs) for efficient computation.
Specifically, we perform the change of variables $\widetilde{\varphi}(x):=\frac{1}{2}\|x\|^{2}-\varphi(x)$ and $\widetilde{\psi}(y):=\frac{1}{2}\|y\|^{2}-\psi(y)$ and then, the constraint in the dual problem can be written as $\widetilde{\varphi}(x)+\widetilde{\psi}(y)\geq xy$. Moreover, it is not too difficult to show that $\widetilde{\varphi}$ is convex and $\frac{1}{2}\|y\|^{2}-\psi(y)=\frac{1}{2}\|y\|^{2}-\varphi^{c}(y)=\widetilde{\varphi}^{*}(y)$ is the Legendre transform of $\widetilde{\varphi}$ at optimality, see Proposition 1.21 in [43]. Then the optimization problem (3.8) can be expressed as
(5.5) $\sup_{\widetilde{\varphi}}\left\{\mathbb{E}_{P}[-\widetilde{\varphi}]-\epsilon\log\mathbb{E}_{Q}[e^{\widetilde{\varphi}^{*}/\epsilon}e^{-\|y\|^{2}/2\epsilon}]\right\}+C=-\inf_{\widetilde{\varphi}}\left\{\mathbb{E}_{P}[\widetilde{\varphi}]+\epsilon\log\mathbb{E}_{Q}[e^{\widetilde{\varphi}^{*}/\epsilon}e^{-\|y\|^{2}/2\epsilon}]\right\}+C,$
where $C=\frac{1}{2}\mathbb{E}_{P}\|x\|^{2}$ and $\widetilde{\varphi}$ is a convex and continuous function on $X$. This reformulation can also be interpreted as expressing the objective in terms of a modified cost function $c(x,y)=-xy$ and a perturbed measure $Q$ obtained by multiplying $Q$ by a standard Gaussian density with variance $\epsilon$ followed by re-normalization.
In practice, we solve this by parameterizing $\widetilde{\varphi}$ with an input-convex neural network in the class ICNN. From the proof of Theorem 3.3 in [27] we have that
$\langle y,\nabla g(y)\rangle-\widetilde{\varphi}(\nabla g(y))\leq\langle y,\nabla\widetilde{\varphi}^{*}(y)\rangle-\widetilde{\varphi}(\nabla\widetilde{\varphi}^{*}(y))=\widetilde{\varphi}^{*}(y),$
for all differentiable $g,\widetilde{\varphi}$. Moreover, the supremum of the left-hand-side over convex functions $g\in\texttt{CVX}$ achieves the Legendre transform $g=\widetilde{\varphi}^{*}$. Given that $x\mapsto\epsilon\log\mathbb{E}_{Q}[e^{x/\epsilon}e^{-\|y\|^{2}/2\epsilon}]$ is an increasing function, we have that
$\sup_{g\in\texttt{CVX}}\epsilon\log\mathbb{E}_{Q}[e^{(\langle y,\nabla g(y)\rangle-\widetilde{\varphi}(\nabla g(y)))/\epsilon}e^{-\|y\|^{2}/2\epsilon}]\leq\epsilon\log\mathbb{E}_{Q}[e^{-\widetilde{\varphi}^{*}(y)}e^{-\|y\|^{2}/2\epsilon}].$
Using this result, we can compute the OT divergence in (5.5) by solving the variational problem
(5.6) $\inf_{\widetilde{\varphi}\in\texttt{ICNN}}\sup_{g\in\texttt{ICNN}}\left\{\mathbb{E}_{P}[\widetilde{\varphi}]+\epsilon\log\mathbb{E}_{Q}[e^{(\langle y,\nabla g(y)\rangle-\widetilde{\varphi}(\nabla g(y)))/\epsilon}e^{-\|y\|^{2}/2\epsilon}]\right\}.$
5.3. Generative flow-based algorithms
Here, we demonstrate that a specific class of normalizing flows, known as the optimal-transport flow (OT-flow) introduced in [34], is derived from a proximal OT divergence. The OT-flow is a type of generative model that was formulated as a stable version of the normalizing flows [40, 48, 47]. Given only samples from the distribution $Q$, a normalizing flow is used to learn a transformation that maps these samples to a simpler, typically known distribution, such as a standard Gaussian. The learned transformation is called the flow. The inverse of this flow serves as the generative model, which transforms samples from the simpler distribution $P$ into new samples that approximate the complex distribution $Q$. In [34], the flow is found via maximum likelihood estimation, which is regularized by the Wasserstein distance and the connection with the proximal OT divergences is immediate:
$\min_{v,\rho}\Bigg{\{}\mathbb{E}_{\rho(x,T)}\left[\log\frac{\rho(x,T)}{P(x)}\right]+\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{1}{2}|v(x,t)|^{2}\rho(x,t)dx\,dt$
$\hskip 56.9055pt:\frac{\partial\rho}{\partial t}+\nabla\cdot(v\rho)=0,\,\rho(x,0)=Q(x)\Bigg{\}}$
(5.7) $=2T\inf_{R}\left[\mathfrak{D}_{\mathrm{KL}}(R\|P)+\frac{1}{2T}W_{2}^{2}(Q,R)\right]=\mathfrak{D}^{2}_{\mathrm{KL},2T}(Q\|P)$
This framework allows us to compute our divergences and the and proximal minimizer when KL and Wasserstein-2 efficiently by Neural ODEs [11] and OT-flow implemented by the tools outlined in Section 3 of [34]. Training an OT-flow requires both forward and backward simulations of invertible flows, and can be unstable in high-dimensional examples where data often lies on a lower-dimensional manifold. In fact, the choice of KL in (5.3) does not guarantee the existence of an optimizer due to the lack of absolute continuity, and there is no invertible mapping between the reference distribution being often a $d$-dimensional Gaussian and the target distribution supported on a lower-dimensional manifold. $W_{1}\bigoplus W_{2}$ proximal generative flows in [21], address these challenges, as they do not require the inversion of the flow or absolute continuity between the measures.
1. Use OT-flow to compute the intermediate measure (proximal).
6. Gradient flows for proximal OT divergences and connections to Probability flows
In this section, we consider $Q$ as a fixed target measure-typically arising from observed data-and explore how the proximal OT divergence $\mathfrak{D}^{c}_{\varepsilon}(P\|Q)$, can be leveraged to transport a source measure $P$ toward $Q$ by building gradient flows for $\mathfrak{D}^{c}_{\varepsilon}(P\|Q)$. This has been extensively studied in the recent work [20] for the special case of the 1-Wasserstein proximal divergence under the assumption that $Q$ has a finite first moment, and [13] where the authors demonstrate the robustness of the 1-Wasserstein $\alpha$-divergences in generative modeling without requiring any assumptions on $Q$. Finally, we establish connections with probability flow ODEs. These ODEs provide a deterministic framework for evolving distributions over time and have recently emerged as powerful tools in generative modeling.
A central step in constructing the gradient flow is the computation of the variational derivative of the divergence, denoted informally as $\frac{\delta\mathfrak{D}^{c}_{\varepsilon}(P\|Q)}{\delta P}$ given in Theorem 4.1, not only facilitates the definition of the flow but also sheds light on the regularizing influence of the entropic term. In particular, the entropic regularization ensures the existence and uniqueness of a well-defined variational derivative–something that is generally lacking in the classical optimal transport setting (see, e.g., Section 7.2 of [41]). Conversely, adding the optimal transport term to the KL divergence alters its properties in a meaningful way, given that the KL divergence itself already has a well-defined variational derivative, i.e., $\frac{\delta\mathfrak{D}_{\mathrm{KL}}(P\|Q)}{\delta P}=1+\log\frac{dP}{dQ}$ provided that $P$ is absolutely continuous to $Q$. Regularizing with the optimal transport, the variational derivative can be extended to apply to more general, non-absolutely continuous measures $P$. Let $P$ be a source probability distribution and $Q$ an arbitrary target distribution. At the formal level, a gradient flow on the space of probability measures obtained as
(6.1) $\partial_{t}P_{t}={\rm div}\left(P_{t}\nabla\frac{\delta\mathfrak{D}^{c}_{\varepsilon}(P_{t}\|Q)}{\delta P_{t}}\right)\,,\quad P_{0}=P$
As noted in Remark 2.7, the gradient flow (2.14) in the JKO scheme shares the same structure as (6.1) and finds the approximate solution at each time-step by solving (2.15). Here, we aim to leverage the formulation in (6.1) to construct generative particle algorithms as discussed next. Recalling the variational derivative of the proximal OT-divergence with the KL divergence given in Theorem 4.1, i.e.,
$\frac{\delta\mathfrak{D}_{\mathrm{KL},\varepsilon}^{c}(P\|Q)}{\delta P}=\widehat{\phi}^{*}\quad\mathrm{with}\quad\widehat{\phi}^{*}(x)=\inf_{y\in\textrm{supp}(Q)}\left\{c(x,y)-\psi^{*}(y)\right\}\,,$
and
$(\phi^{*},\psi^{*})=\textrm{argmax}_{(\phi,\psi)\in\Phi_{c}}\left\{\mathbb{E}_{P}[\phi]-\varepsilon\log\mathbb{E}_{Q}[e^{-\psi/\varepsilon}]\right\}$
we construct the gradient flows for proximal OT divergences as follows
$\partial_{t}P_{t}$ $={\rm div}\left(P_{t}\nabla\widehat{\phi}_{t}^{*}\right)\,,\quad P_{0}=P,$
(6.2) $\widehat{\phi}^{*}_{t}(x)$ $=\inf_{y\in\textrm{supp}(Q)}\left\{c(x,y)-\psi^{*}_{t}(y)\right\}$
$(\phi_{t}^{*},\psi_{t}^{*})$ $=\textrm{argmax}_{(\phi,\psi)\in\Phi_{c}}\left\{\mathbb{E}_{P_{t}}[\phi]-\varepsilon\log\mathbb{E}_{Q}[e^{-\psi/\varepsilon}]\right\}$
In the context of generative models, both the target distribution $Q$ and the generative model $P_{t}$ in (6.1) are accessible only through samples and their corresponding empirical distributions. It has been mentioned that $\mathfrak{D}_{\mathrm{KL},\varepsilon}^{c}(P\|Q)$ can directly compare such singular distributions as ensured by the upper bound in (2.10). We now turn to the Lagrangian formulation of the PDE (6.1), which corresponds to its underlying ODE/variational representation.
$\frac{dY_{t}}{dt}$ $=-\nabla\widehat{\phi}_{t}^{*}(Y_{t})\,,\quad Y_{0}\sim P_{0}=P$
(6.3) $\widehat{\phi}^{*}_{t}(x)$ $=\inf_{y\in\textrm{supp}(Q)}\left\{c(x,y)-\psi^{*}_{t}(y)\right\}$
$(\phi_{t}^{*},\psi_{t}^{*})$ $=\textrm{argmax}_{(\phi,\psi)\in\Phi_{c}}\left\{\mathbb{E}_{P_{t}}[\phi]-\varepsilon\log\mathbb{E}_{Q}[e^{-\psi/\varepsilon}]\right\}$
To construct a particle algorithm based on (6), we begin with the input data $(X^{(i)})_{i=1}^{N}$ from the unknown target distribution $Q$ and $(Y^{(j)})_{j=1}^{M}$ drawn from a source distribution $P$. The corresponding empirical measures are $\hat{Q}^{N}$ and $\hat{P}_{0}^{M}$. The Euler time discretization of (6) takes the form:
$Y_{n+1}^{(j)}$ $=Y_{n}^{(j)}-\Delta t\nabla\widehat{\phi}_{n}^{*}(Y_{n}^{(j)})\,,\quad Y_{0}^{(j)}=Y^{(j)}\sim P_{0},\qquad j=1,\dots,M$
$\phi^{*}_{n}(Y_{n}(j))$ $=\min_{i=1,\dots,N}\left\{c(Y_{n}(j),X^{(i)})-\psi^{*}_{n}(X^{(i)})\right\},\qquad j=1,\dots,M$
(6.4) $(\phi_{n}^{*},\psi_{n}^{*})$ $=\textrm{argmax}_{(\phi,\psi)\in\Phi_{c}}\left\{\frac{\sum_{j=1}^{M}\phi(Y_{n}^{(i)})}{N}-\varepsilon\log\frac{\sum_{i=1}^{N}e^{-\psi(X^{(i)})/\varepsilon}}{N}\right\}$
where we remind that $\Phi_{c}=\left\{(\phi,\psi)\in C_{b}(\mathbb{R}^{d}\times\mathbb{R}^{d}),\phi\oplus\psi\leq c\right\}$. The discrete scheme (6) can be interpreted as a generative algorithm that transports the initial source samples $(Y_{0}^{(j)})_{j=1}^{M}$ from $P_{0}$ (typically a distribution that is easy to sample) over a time horizon $T=n_{T}\Delta t$, where $n_{T}$ is the total number of steps, to a new set of generated data $(Y_{n_{T}}^{(j)})_{j=1}^{M}$ that approximate samples from $Q$.
In [20], the particular case of $\mathfrak{D}_{\mathrm{KL},\varepsilon}^{1}$–the proximal OT divergence with Wasserstein-1 distance–is studied. Consequently, the optimization problem in (6) is formulated over the space of Lipschitz functions. To make this tractable in practice, this function space $\mathrm{Lip_{\frac{1}{\varepsilon}}({\mathbb{R}}^{d})}$ is approximated using neural networks $\Gamma^{NN}_{\frac{1}{\varepsilon}}$, i.e., $\phi_{n}^{*}=\textrm{argmax}_{\phi\in\Gamma^{NN}_{\frac{1}{\varepsilon}}}\left\{\frac{\sum_{j=1}^{M}\phi(Y_{n}^{(i)})}{N}-\varepsilon\log\frac{\sum_{i=1}^{N}e^{\phi(X^{(i)})/\varepsilon}}{N}\right\}$ where the Lipschitz constraint is enforced through spectral normalization techniques [29]. Moreover, the optimization problem in the discrete scheme (6) for the gradient flow of $\mathfrak{D}_{\mathrm{KL},\varepsilon}^{2}$ can be implemented by solving the variational problem (5.6) using input convex neural networks.
Connections to Probability Flows and score-matching: Next, we explore the connection between gradient flows and probability flow ODEs (i.e., deterministic sampling), [28]. Given an initial condition $p_{0}$, the Fokker-Planck equation given by
$\partial_{t}p(t,y)=-{\rm{div}}(f(t,y)p(t,y))+\frac{1}{2}g^{2}(t)\Delta p(t,y)$
describes the evolution of $p_{t}(y)$ for the random variable $Y_{t}$ following the forward SDE
$dY_{t}=f(t,Y_{t})dt+g(t)dW_{t}$
where $f(t,\cdot):\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $g(t)\in\mathbb{R}$ are the drift and diffusion coefficient respectively, and $W_{t}$ is a standard Brownian motion. The Laplacian of $p$ can be written as a transport term driven by the score function $\nabla\log p$, i.e.,
$\Delta p={\rm{div}}(p\nabla\log p)$
and the Fokker-Planck equation is written as
$\frac{\partial p_{t}(y)}{\partial t}=-{\rm div}\left(p_{t}(y)v(t,y)\right),\qquad v(t,y)=f(t,y)-\frac{g(t)^{2}}{2}\underbrace{\nabla\log p_{t}(y)}_{\textrm{score function of $p_{t}$}}$
The corresponding probability flow ODE is given by
$\frac{dY_{t}}{dt}=v(t,Y_{t}),\quad Y_{0}\sim p_{0}$
This perspective also forms the core idea in [46], which used as a deterministic alternative to the reverse-time stochastic differential equation (SDE), offering a new way to generate samples and compute exact likelihoods. In fact, the authors first use the forward SDE to gradually add noise to the data, then train a time-dependent score-based model $s_{\theta}(y)$ by minimizing the Fisher divergence between the model and the data distributions defined as
(6.5) $\mathbb{E}_{p(y)}\left[\|s_{\theta}(y)-\nabla\log p_{t}(y)\|^{2}_{2}\right]$
using score matching techniques (e.g., [45]), and finally generate samples by solving the corresponding probability flow ODE.
In our gradient flow perspective, starting from the simple case of the classical KL-divergence and computing its variational derivative, reveals a key connection to score-based modeling. Specifically, the gradient of the variational derivative of the KL divergence–equal to the gradient of the discriminator–can be written in terms of the score function: $\nabla\frac{\delta\mathfrak{D}_{\mathrm{KL}}(p_{t}\|q)}{\delta p_{t}}=\nabla\phi^{*}_{t}=\nabla\log\frac{p_{t}}{q}=\nabla\log p_{t}-\nabla\log q$ and thus
(6.6) score function of pt =$\textrm{score function of $p_{t}$ }=\nabla\log p_{t}=\nabla\phi^{*}_{t}+\nabla\log q$
This formulation illustrates how score-based modeling naturally arises from the KL-divergence, thereby linking our gradient flow framework to the probability flow ODE. In fact, for specific choices of the functions $g(t)$ and $f(t,y)$, this yields the gradient flow of the KL divergence, which is further generalized by the gradient flow associated with proximal OT divergences in (6.1). From a computational standpoint, rather than training a score-based model directly, we solve the variational problem defined by (6) using convex optimization techniques, as detailed in Section 5.1 and neural networks and discriminators in analogy to GANs, [20].